OpenStudy (anonymous):

Solving equations: 2/x-3=8

OpenStudy (shane_b):

Is it \[\frac{2}{x-3}=8\]or \[\frac{2}{x}-3=8\]

OpenStudy (anonymous):

the first one

OpenStudy (shane_b):

Ok, multiply both sides by x-3 to get rid of the fraction bar. What does that leave you with?

OpenStudy (anonymous):

2=8(x-3)?

OpenStudy (shane_b):

Exactly...now distribute the 8...

OpenStudy (anonymous):

2=8x-24 which goes to 26=8x divide both sides by 8 x=3.25?

OpenStudy (shane_b):

That's correct :)

OpenStudy (anonymous):

Alright thanks :)

OpenStudy (shane_b):

you're welcome :)

OpenStudy (shane_b):

Easier maybe...but that answer is wrong :)

OpenStudy (anonymous):

Argh! How silly of me. =P